Question 10 1 pts The molar solubility (in M) of PbCl2 in a 0.20 M solution of HCl is This molar solubility is...than the molar solubility of PbCl2 in pure water. The Ksp of PbCl2 is 1.6x10 O 4.0x10-4 M, higher O 4.0x104 M, lower O 80x105 M, higher 8.0 x 105 M, lower
Answer #1

PbCl2 ===> Pb2+ + 2Cl-
[Pb2+] * [Cl-]2 = 1.6*10-5
x * (0.2+x)2 =1.6*10-5

x*0.04 = 1.6*10-5

it will be B